# Benchmarking Unsupervised Online IDS for Masquerade Attacks in CAN

### v2

@@ IV-B2 Testing Phase (P2)
In the testing stage, data is processed using a sliding window approach (see more details in Section IV-A). Specifically, smaller sections of $\mathbf{X}_{test}$ of length $\omega$ are selected to create a view of the steaming ($\in\mathbb{R}^{\omega\times n}$). Benchmarked methods consume this batch of data to arrive at conclusions.
2) Model Fitting (T2):
We fit the benhmarked methods with their specific models on $\mathbf{X}_{test}$. The specifics of the fitted models are discussed in Section IV-C.
@@
This step entails a data-driven decision-making for deciding if a particular window is containing an anomaly or not. The specifics of the criteria for detection are discussed in Section IV-C.
IV-C Benchmarked Algorithms
This subsection describes in detail each of the four algorithms used in this research. Note that we focus on non DL-based algorithms because the dynamic nature of CAN data makes it hard for DL-based methods to extract useful features, which reduces their performance [24]. For the sake of readability, we associate each algorithm name of the first author and the last two digits of the publication year (when applicable). Specifically, we describe the following methods: Matrix Correlation Distribution (see Section IV-C1), Matrix Correlation Correlation (see Section IV-C2), Ganesan17 [23] (see Section IV-C3), and Moriano22 [15] (see Section IV-C4).
We notice that none of the benchmarked algorithms has been distributed with a corresponding implementation. In addition, many of their implementation details are not provided. To address these limitations, we make available our referenced implementations and provide a detailed description of the assumptions and parameters tested in each of the algorithms. This allows practicioners in the area the opportunity to easily replicate our experiments and benchmark novel signal-based detection algorithms with respect to the state of the art. Note that becnhmarked algorithms produce anomaly scores in the range 0 to 1 that is being used for computing evaluation metrics.
All method computes pairwise Pearson correlations among time series in $\mathbf{X}_{train}$ (and $\mathbf{X}_{test}$) obtaining $\mathbf{R}_{train}\in\mathbb{R}^{n\times n}$ (and corresponding $\mathbf{R}_{test}\in\mathbb{R}^{n\times n}$), symmetric matrices whose $r_{ij}$ entry is the Pearson correlation coefficient [54] between the signals $s_{i}$ and $s_{j}$. Pearson correlation values that are close to ±$\pm 1.0$ indicate strong positive or negative correlation. Note that time series that have positive correlations are expected to move jointly (i.e., when one time series increases or decreases, the other time series also increases or decreases). As vehicle dynamics are constrained by laws of nature, we expect that: (1) clusters of correlated signals emerge, for example, as there is an increase in the speed of the vehicle, there is a subsequent increase in the speedometer reading and the speed of the four wheels, and (2) such clusters are disturbed upon a cyberattack given that time series correlations could be broken or changed significantly.
@@ IV-C Benchmarked Algorithms
Note that the benchmarked algorithms suit streaming environments. They process small data batches and do not rely on large, representative training datasets, unlike DL-based methods.
IV-C1 Implementation of Matrix Correlation Distribution
@@
IV-C2 Implementation of Matrix Correlation Correlation
This method considers the upper triangular entries of $\mathbf{U}_{train},\mathbf{U}_{test}$ (the signals’ pairwise Pearson correlation coefficients) and computes the statistical significance of their Spearman correlation [56], a measure in $[-1,1]$ of their members ordinal rank. Simply unravel the upper triangular entries of $\mathbf{U}_{train},\mathbf{U}_{test}$ into two sequences and compute their Spearman correlation. Note that Spearman correlation furnishes, alongside the correlation value, a $p$-value giving the likelihood that the two input vectors are from uncorrelated systems (so low $p$-value corresponds to the two input vectors coming from correlated systems). We focus on the $p$-value of the Spearman correlation as this captures the reliability of the correlation. The Spearman correlation coefficient is better suited for non-normally distributed continuous data, ordinal data, or data with potential outliers as a measure of monotonic association [57]. Pearson correlation is also scaled in the range ±$\pm 1$ indicating a constantly increasing or decreasing monotonic association while 00 indicating no monotonic association. The Spearman correlation null hypothesis is that there is no correlation between the strictly upper triangular correlations in training and testing (opposite to the first test above); hence, we use the $p$-value as the probability score. Here, higher $p$-values suggest not rejecting the null hypothesis; this translates to a higher anomaly score. Conversely, lower $p$-values suggest rejecting the null hypothesis; this translates to a lower anomaly score.
IV-C3 Implementation of Ganesan17
@@ IV-C4 Implementation of Moriano22
For detection, in each time window of length $\omega$, we compute the clustering similarity $s$ between the hierarchical clustering from training (computed once) and the hierarchical clustering computed during that specific time window. Here, similarity is used as a proxy of the distance between both hierarchical clusterings. We compute the similarity between hierarchical clustering using the CluSim method [61] based on the work by Gates et al. [62]. The similarity value provided by this method, $s$, lies in the range $[0,1]$, where 0 implies maximally dissimilar clusters, and 1 corresponds to identical clusters. Therefore, we report the anomaly score (or probability of detection) as $1-s$.
IV-C5 Algorithms’ Hyperparameter Setup
@@
Matrix Correlation Distribution and Matrix Correlation Correlation algorithms do not require hyperparameters. They report anomaly scores based on $p$-values from the Mann-Whitney U test and Spearman correlation, respectively. For Ganesan17, we used the similar values as in Ganesan et al. [23], i.e., we set both $eps$ and $min\_samples$ to 1 to ensure a reasonable similarity between time series to be consider as part of the same cluster and allow individual time series to be part of its own cluster, respectively. For Moriano22, we used the sames values as in Moriano et al. [15], i.e., setting $r=-5.0$ and $\alpha=0.9$. Here, $r$ is a scaling parameter that defines the relative importance of memberships at different levels of the hierarchy. That is, the larger $r$, the more emphasis on comparing lower levels of the dendrogram (zoom in). In addition, $\alpha$ is a parameter that controls the influence of hierarchical clusterings with shared lineages. That is, the larger $\alpha$, the further the process will explore from the focus data element, so more of the cluster structure is taken into account into the comparison.
IV-D Dataset
@@
We ran the proposed benchmark on the ROAD dataset [10]. The ROAD dataset is a publicly available dataset for CAN IDS that contains benign and attack CAN data collected from a real vehicle. An important feature of ROAD is that it provides samples of masquerade attacks which are the types of attacks we are interested in detecting with the bechmarked methods. Note that masquerade attacks in ROAD are identical to the corresponding targeted ID captures but with the ambient frames of the targeted ID removed during the attack period to simulate the masquerade attack. The ROAD dataset also provides already translated CAN time series from log files. By using ROAD in this study, we ensure comparing the performance of proposed methods with realistic, verified, and labelled attacks as opposed to synthetic ones. This allows performing the comparison under realistic conditions.
@@
We used the longest benign capture in the ROAD dataset (i.e., ambient-highway-street-driving-long) for learning the normal state of the system ($\approx$ 1 hour of data). We report results on the proposed benchmark in each of the masquerade attacks provided in the ROAD dataset. Note that each of the masquerade attack files in ROAD contains hundreds of time series coming from hundreds IDs each of them with up to a few dozen signals. Table I shows the specific files we used from ROAD. In particular, we used the following attacks in increasing level of detection difficulty: correlated signal, max speedometer, max engine coolant temperature, reverse light on, and reverse light off. In the correlated signal attack, the correlation of the four wheel speed values is manipulated by changing their individual values. In the max speedometer and max engine coolant attacks, the speedometer and coolant temperature values are changed to their maximum. In the reverse light attacks, the state of the reverse light is modified so that it does not match the current gear of the vehicle (i.e., reverse light is off when the vehicle is in reverse and the reverse light is on when the vehicle is in forward gear).
TABLE I: Masquerade attack characteristics in the ROAD dataset [10].
[…]
IV-E Evaluation Metrics
@@
This study focuses on comparing the detection capability of the benchmarked algorithms on the ROAD dataset. As such, we report summarized results over a range of detection thresholds and focus on the area under the receiver operating characteristic curve (AUC-ROC) metric. The AUC-ROC reflects the overall performance of the benchmarked algorithms based on the variation of the true positive rate (TPR) with respect to the false positive rate (FPR) at various thresholds [63]. Mathematically, it is expressed as
[…]
@@
IV-F Performance Metrics
@@
For assessing the real-time capabilities of the benchmarked algorithms, we also report the latency that these methods have to infer detection decisions on the sliding windows being processed. In doing so, we report the testing time per window (TTW), for each combination of window length ($\omega$) and offset ($\delta$), following the concept of testing time per packet used by Nichelini et al. [20]. We define TTW as
TTW=Total Detection TimeNumber of $\text{TTW}=\frac{\text{Total Detection Time}}{\text{Number of Windows}}$
@@
Our experiments were conducted on a AMD EPYC 7702 64-Core Processor Central Processing Unit (CPU), 256 GB of RAM, running Ubuntu 22.04.4 LTS (Jammy Jellyfish). TTW provides a way to estimate the applicability of the bechmarked algorithms in a real-time setting.
V Results
@@
In this section, we detail the experimental evaluation of the benchmarked algorithms on the ROAD dataset. All methods were implemented in Python 3.8.12. We use the sklearn.metrics module to compute the evaluation metrics and the timeit module to compute the TTW of each benchmarked algorithm. To enable the reproducibility of the benchmarked methods, we make the code available in GitHub. [^1] https://github.com/pmoriano/benchmarking-unsupervised-online-IDS-masquerade-attacks We conduct three different but complementary analysis.
@@
First, we compute and contrast the detection performance (based on AUC-ROC) of each algorithm per attack category in the ROAD dataset. In doing so, we display heatmaps (Section V-A). We hypothesize that each algorithm is better suited for detecting specific attacks in the ROAD dataset and that there is a range of $(\omega,\delta)$ combinations that maximizes AUC-ROC.
@@
Second, we summarize evaluation metrics using summary statistics derived from the heatmaps (Section V-B). Specifically, we report the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), including the combination of window length and offset that produces it, and maximum ($\max$), including the combination of window length and offset that produces it. This analysis supports our hypothesis that there are better suited detection algorithms for each attack category. Overall, Moriano22 detection algorithm produces, on average, the best detection performance among attacks in the ROAD dataset. In contrast, Ganesan17 detection algorithm produces the worst detection performance among attacks in the ROAD dataset. Furthermore, we found that the easiest attacks to detect are the max_engine_coolant_attack and correlated_attack. Furthermore, we find that the max_speedometer_attack and light_on_attack are the hardest to detect. We also optimize the hyperparameters for the best-performing method, i.e., Moriano22. Using grid search, we demonstrate the potential for AUC-ROC improvement through hyperparameter tuning.
@@
Lastly, we summarize TTW as a proxy of performance (Section V-C). In particular, we report the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), including the combination of window length and offset that produces it, and maximum ($\max$), including the combination of window length and offset that produces it. This analysis reveals that more sophisticated algorithms (e.g., Ganesan17 and Moriano22) incur performance overhead for producing better detection evaluation metrics. Overall, Matrix Correlation Distribution produces, on overage, the quickest TTW among attacks in the ROAD dataset. In contrast, the Moriano22 detection algorithm runs at least four times slower, as it uses a standard AHC implementation rather than a faster approximate alternative [65]. It is worth noting that detection algorithms that run faster (i.e., Matrix Correlation Distribution and Matrix Correlation Correlation) perform even better than Ganesan17 (see Table IV). We did not found significant differences on TTW per attack categories.
V-A Evaluation Metrics Contrast

### v3

@@ 4.2.2 Testing Phase (P2)
In the testing stage, data is processed using a sliding window approach (see more details in Section 4.1). Specifically, smaller sections of $\mathbf{X}_{test}$ of length $\omega$ are selected to create a view of the steaming ($\in\mathbb{R}^{\omega\times n}$). Evaluated methods consume this batch of data to arrive at conclusions.
2) Representation Extraction (T2):
During the testing phase (P2), we apply the evaluated methods to extract statistical representations from the windowed test data $\mathbf{X}_{test}$ and compare them to those obtained from the training data (P1). No model fitting is performed on the test set. The evaluated methods are discussed in Section 4.3.
@@
This step entails a data-driven decision-making for deciding if a particular window is containing an anomaly or not. The specifics of the criteria for detection are discussed in Section 4.3.
4.3 Evaluated Algorithms
This subsection describes in detail each of the four algorithms used in this research. Note that we focus on non DL-based algorithms because the dynamic nature of CAN data makes it hard for DL-based methods to extract useful features, which reduces their performance [24]. For the sake of readability, we associate each algorithm name of the first author and the last two digits of the publication year (when applicable). Specifically, we describe the following methods: Matrix Correlation Distribution (see Section 4.3.1), Matrix Correlation Correlation (see Section 4.3.2), Ganesan17 [23] (see Section 4.3.3), and Moriano22 [15] (see Section 4.3.4).
We notice that none of the evaluated algorithms has been distributed with a corresponding implementation. In addition, many of their implementation details are not provided. To address these limitations, we make available our referenced implementations and provide a detailed description of the assumptions and parameters tested in each of the algorithms. This allows practicioners in the area the opportunity to easily replicate our experiments and comparative evaluation novel signal-based detection algorithms with respect to the state of the art. Note that becnhmarked algorithms produce anomaly scores in the range 0 to 1 that is being used for computing evaluation metrics.
All method computes pairwise Pearson correlations among time series in $\mathbf{X}_{train}$ (and $\mathbf{X}_{test}$) obtaining $\mathbf{R}_{train}\in\mathbb{R}^{n\times n}$ (and corresponding $\mathbf{R}_{test}\in\mathbb{R}^{n\times n}$), symmetric matrices whose $r_{ij}$ entry is the Pearson correlation coefficient [59] between the signals $s_{i}$ and $s_{j}$. Pearson correlation values that are close to $\pm 1.0$ indicate strong positive or negative correlation. Note that time series that have positive correlations are expected to move jointly (i.e., when one time series increases or decreases, the other time series also increases or decreases). As vehicle dynamics are constrained by laws of nature, we expect that: (1) clusters of correlated signals emerge, for example, as there is an increase in the speed of the vehicle, there is a subsequent increase in the speedometer reading and the speed of the four wheels, and (2) such clusters are disturbed upon a cyberattack given that time series correlations could be broken or changed significantly.
@@ 4.3 Evaluated Algorithms
Note that the evaluated algorithms suit streaming environments. They process small data batches and do not rely on large, representative training datasets, unlike DL-based methods.
4.3.1 Implementation of Matrix Correlation Distribution
@@
4.3.2 Implementation of Matrix Correlation Correlation
This method considers the upper triangular entries of $\mathbf{U}_{train},\mathbf{U}_{test}$ (the signals’ pairwise Pearson correlation coefficients) and computes the statistical significance of their Spearman correlation [61], a measure in [1,1][-1,1] of their members ordinal rank. Simply unravel the upper triangular entries of $\mathbf{U}_{train},\mathbf{U}_{test}$ into two sequences and compute their Spearman correlation. Note that Spearman correlation furnishes, alongside the correlation value, a $p$-value giving the likelihood that the two input vectors are from uncorrelated systems (so low $p$-value corresponds to the two input vectors coming from correlated systems). We focus on the $p$-value of the Spearman correlation as this captures the reliability of the correlation. The Spearman correlation coefficient is better suited for non-normally distributed continuous data, ordinal data, or data with potential outliers as a measure of monotonic association [62]. Pearson correlation is also scaled in the range $\pm 1$ indicating a constantly increasing or decreasing monotonic association while 0 indicating no monotonic association. The Spearman correlation null hypothesis is that there is no correlation between the strictly upper triangular correlations in training and testing (opposite to the first test above); hence, we use the $p$-value as the probability score. Here, higher $p$-values suggest not rejecting the null hypothesis; this translates to a higher anomaly score. Conversely, lower $p$-values suggest rejecting the null hypothesis; this translates to a lower anomaly score.
4.3.3 Implementation of Ganesan17
@@ 4.3.4 Implementation of Moriano22
For detection, in each time window of length $\omega$, we compute the clustering similarity $s$ between the hierarchical clustering from training (computed once) and the hierarchical clustering computed during that specific time window. Here, similarity is used as a proxy of the distance between both hierarchical clusterings. We compute the similarity between hierarchical clustering using the CluSim method [66] based on the work by Gates et al. [67]. The similarity value provided by this method, $s$, lies in the range $[0,1]$, where 0 implies maximally dissimilar clusters, and 1 corresponds to identical clusters. Therefore, we report the anomaly score (or probability of detection) as 1s1-s.
4.3.5 Algorithms’ Hyperparameter Setup
@@
Matrix Correlation Distribution and Matrix Correlation Correlation algorithms do not require hyperparameters. They report anomaly scores based on $p$-values from the Mann-Whitney U test and Spearman correlation, respectively. For Ganesan17, we used the similar values as in Ganesan et al. [23], i.e., we set both epseps and min_samplesmin\_samples to 1 to ensure a reasonable similarity between time series to be consider as part of the same cluster and allow individual time series to be part of its own cluster, respectively. For Moriano22, we used the sames values as in Moriano et al. [15], i.e., setting r=5.0r=-5.0 and $\alpha=0.9$. Here, $r$ is a scaling parameter that defines the relative importance of memberships at different levels of the hierarchy. That is, the larger $r$, the more emphasis on comparing lower levels of the dendrogram (zoom in). In addition, $\alpha$ is a parameter that controls the influence of hierarchical clusterings with shared lineages. That is, the larger $\alpha$, the further the process will explore from the focus data element, so more of the cluster structure is taken into account into the comparison.
4.4 Dataset
@@
We ran the proposed comparative evaluation on the ROAD dataset [10]. The ROAD dataset is a publicly available dataset for CAN IDS that contains benign and attack CAN data collected from a real vehicle. An important feature of ROAD is that it provides samples of masquerade attacks which are the types of attacks we are interested in detecting with the evaluated methods. Note that masquerade attacks in ROAD are identical to the corresponding targeted ID captures but with the ambient frames of the targeted ID removed during the attack period to simulate the masquerade attack. The ROAD dataset also provides already translated CAN time series from log files. By using ROAD in this study, we ensure comparing the performance of proposed methods with realistic, verified, and labelled attacks as opposed to synthetic ones. This allows performing the comparison under realistic conditions.
@@
We used the longest benign capture in the ROAD dataset (i.e., ambient-highway-street-driving-long) for learning the normal state of the system ($\approx$ 1 hour of data). We report results on the proposed comparative evaluation in each of the masquerade attacks provided in the ROAD dataset. Note that each of the masquerade attack files in ROAD contains hundreds of time series coming from hundreds IDs each of them with up to a few dozen signals. Table 1 shows the specific files we used from ROAD. In particular, we used the following attacks in increasing level of detection difficulty: correlated signal, max speedometer, max engine coolant temperature, reverse light on, and reverse light off. In the correlated signal attack, the correlation of the four wheel speed values is manipulated by changing their individual values. In the max speedometer and max engine coolant attacks, the speedometer and coolant temperature values are changed to their maximum. In the reverse light attacks, the state of the reverse light is modified so that it does not match the current gear of the vehicle (i.e., reverse light is off when the vehicle is in reverse and the reverse light is on when the vehicle is in forward gear).
Table 1: Masquerade attack characteristics in the ROAD dataset [10].
Attack Name Duration (s) Injection Interval (s) Description correlated_signal_1 33.10 [9.19, 30.05] Injects varying values for wheel speeds, causing the vehicle to halt. correlated_signal_2 28.23 [6.83, 28.23] correlated_signal_3 16.96 [4.32, 16.96] max_engine_coolant 25.88 [19.98, 24.17] Injects the maximum value, triggering the coolant warning light. max_speedometer_1 88.02 [42.01, 66.45] Injects the maximum value to be displayed on the speedometer. max_speedometer_2 59.70 [16.01, 47.41] max_speedometer_3 86.77 [9.52, 70.59] reverse_light_off_1 28.11 [16.63, 23.35] Toggles the reverse light irrespective of the actual gear position. reverse_light_off_2 40.67 [13.17, 36.88] reverse_light_off_3 57.88 [16.52, 40.86] reverse_light_on_1 54.85 [18.93, 38.84] Toggles the reverse light irrespective of the actual gear position. reverse_light_on_2 72.02 [20.41, 57.30] reverse_light_on_3 64.26 [23.07, 46.58]
4.5 Evaluation Metrics
@@
This study focuses on comparing the detection capability of the evaluated algorithms on the ROAD dataset. As such, we report summarized results over a range of detection thresholds and focus on the area under the receiver operating characteristic curve (AUC-ROC) metric. The AUC-ROC reflects the overall performance of the evaluated algorithms based on the variation of the true positive rate (TPR) with respect to the false positive rate (FPR) at various thresholds [68]. Mathematically, it is expressed as
$\text{AUC-ROC}=\int_{0}^{1}\text{TPR}(\text{FPR}^{-1}(x))dx.$
@@
While alternative classification metrics (e.g., F1-score, MCC) are valuable in fixed-threshold offline evaluations, we intentionally focus on metrics best aligned with streaming contexts. Our public codebase allows easy computation of these additional metrics using preferred thresholding strategies.
4.6 Performance Metrics
@@
For assessing the real-time capabilities of the evaluated algorithms, we also report the latency that these methods have to infer detection decisions on the sliding windows being processed. In doing so, we report the testing time per window (TTW), for each combination of window length ($\omega$) and offset ($\delta$), following the concept of testing time per packet used by Nichelini et al. [20]. We define TTW as
$\text{TTW}=\frac{\text{Total Detection Time}}{\text{Number of Windows}}$
@@
Our experiments were conducted on a AMD EPYC 7702 64-Core Processor Central Processing Unit (CPU), 256 GB of RAM, running Ubuntu 22.04.4 LTS (Jammy Jellyfish). TTW provides a way to estimate the applicability of the evaluated algorithms in a real-time setting.
5 Results
@@
In this section, we detail the experimental evaluation of the evaluated algorithms on the ROAD dataset. All methods were implemented in Python 3.8.12. We use the sklearn.metrics module to compute the evaluation metrics and the timeit module to compute the TTW of each evaluated algorithm. To enable the reproducibility of the evaluated methods, we make the code available in GitHub. [^2] https://github.com/pmoriano/benchmarking-unsupervised-online-IDS-masquerade-attacks We conduct three different but complementary analysis.
@@
First, we compute and contrast the detection performance (based on AUC-ROC) of each algorithm per attack category in the ROAD dataset. In doing so, we display heatmaps (Section 5.1). We hypothesize that each algorithm is better suited for detecting specific attacks in the ROAD dataset and that there is a range of $(\omega,\delta)$ combinations that maximizes AUC-ROC.
@@
Second, we summarize evaluation metrics using summary statistics derived from the heatmaps (Section 5.2). Specifically, we report the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), including the combination of window length and offset that produces it, and maximum ($\max$), including the combination of window length and offset that produces it. This analysis supports our hypothesis that there are better suited detection algorithms for each attack category. Overall, Moriano22 detection algorithm produces, on average, the best detection performance among attacks in the ROAD dataset. In contrast, Ganesan17 detection algorithm produces the worst detection performance among attacks in the ROAD dataset. Furthermore, we found that the easiest attacks to detect are the max_engine_coolant_attack and correlated_attack. Furthermore, we find that the max_speedometer_attack and light_on_attack are the hardest to detect. We also optimize the hyperparameters for the best-performing method, i.e., Moriano22. Using grid search, we demonstrate the potential for AUC-ROC improvement through hyperparameter tuning.
@@
Lastly, we summarize TTW as a proxy of performance (Section 5.3). In particular, we report the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), including the combination of window length and offset that produces it, and maximum ($\max$), including the combination of window length and offset that produces it. This analysis reveals that more sophisticated algorithms (e.g., Ganesan17 and Moriano22) incur performance overhead for producing better detection evaluation metrics. Overall, Matrix Correlation Distribution produces, on overage, the quickest TTW among attacks in the ROAD dataset. In contrast, the Moriano22 detection algorithm runs at least four times slower, as it uses a standard AHC implementation rather than a faster approximate alternative [70]. It is worth noting that detection algorithms that run faster (i.e., Matrix Correlation Distribution and Matrix Correlation Correlation) perform even better than Ganesan17 (see Table 4). We did not found significant differences on TTW per attack categories.
5.1 Evaluation Metrics Contrast
To evaluate how detection metrics are influenced by sliding window configurations, we conducted a fine-grained parameter sweep over key window parameters (e.g., window size $\omega$ and offset $\delta$). These detailed visualizations are important for diagnosing parameter sensitivity and robustness, which are critical in online streaming contexts.